# Stepwise Regression and Pre-trained Edge for Robust Stereo Matching

### v2

@@ III-B5 Disparity Clip-Balanced Weight
In SR-Stereo, the regression disparity is divided into multiple clips with the same range to mitigate the distributional difference among different domains. Ideally, a regression disparity of size $L$ is split into a number of disparity clips of size $1.5m$ and a small disparity clip of size $n$ as follows:
@@ III-B6 Loss Function
We use Smooth L1 loss to supervise the initial disparity and the disparity clips dk because it is more sensitive to large errors:
@@ III-B7 Generalization of the stepwise regression architecture
We observe that although the core idea is different, the stepwise regression architecture is similar to existing iteration-based methods in terms of processing flow, and thus can be easily generalized based on existing methods. In Section x, we also apply the proposed stepwise regression architecture to the classic iteration-based method RAFT-Stereo, and the experimental results also show that the stepwise regression architecture performs better.
@@ III-C1 Edge Estimator
For the proposed DAPE, the reliability of edge pseudo-label is vital to the performance of the fine-tuned model. Existing edge estimation methods usually take only an RGB image as input and use a series of complex 2D CNNs to estimate the contour of object by extracting high-level features. However, this approach significantly increases the computational cost and limits the generalization performance due to the large difference in object classes among different domains.
Figure 5: Architecture of the edge estimator. Compared to edge estimation networks [38, 39, 40] based only on RGB image, our proposed edge estimator is very lightweight.
In this paper, we claim that using disparity for edge estimation is a better and more robust approach because disparity estimation focuses on local geometric information, whose representations are similar across domains. Moreover, compared to RGB image, disparity contains almost no object texture information, which greatly reduces the difficulty of edge estimation. Thus, a lightweight edge estimator is proposed to achieve accurate and robust performance by introducing disparity. The architecture of the edge estimator is shown in Fig. 5. First, a residual block is used to extract the edge feature in the disparity:
@@ III-C2 Edge Map Background Pseudo-label Generation for Target Domain
The proposed edge estimator achieves edge estimation by introducing disparity, and therefore is affected by the accuracy of the disparity. As shown in Fig. 6, existing disparity estimation methods perform badly in reflection regions, which leads to erroneous edges. In this paper, an edge map background pseudo-label is proposed to generate reliable supervision for fine-tuning. Intuitively, in an ideal edge map, both the edge region and non-edge region can be leveraged as supervision to improve the performance of the fine-tuned disparity estimation model. For instance, the disparity map typically exhibits high gradients in edge region, while it remains relatively flat in non-edge region. Erroneous edges in reflection regions do not affect the reliability of the pixels in the non-edge region. Therefore, we can simply utilize pixels in non-edge region as the pseudo-label (i.e., the edge map background pseudo-label):
@@ IV-A Datasets and Evaluation Metrics
Middlebury 2014 [30] consists of two batches of indoor image pairs. The first batch provides 15 training pairs and 15 test pairs, with each scene provided in three different resolutions. The second batch provides 13 additional training pairs, but contains only one resolution. All training pairs are provided with dense hand-labeled ground truth disparity, which ranges from 0 to 300. To evaluate the generalization performance of SR-Stereo, we directly use the training pairs from the first batch. For DAPE, we utilize the additional 13 image pairs to fine-tune the model and evaluate its performance using the 15 training pairs from the first batch. The evaluation metric used is the 2-pixel error.
ETH3D [50] is a grayscale image dataset that includes a variety of indoor and outdoor scenes. It consists of 27 training pairs and 20 test pairs. The dataset provides sparsely labeled ground truth disparities for the training pairs, ranging from 0 to 60 (the smallest among several datasets). Similar to the usage in Middbury 2014, we employ the training pairs to directly assess the generalization performance of the proposed method, using the 1-pixel error as the evaluation metric. Additionally, we use 14 training pairs as fine-tuning samples and 13 training pairs as test samples to evaluate the effectiveness of the proposed Domain Adaptation Based on Pre-trained Edge (DAPE).
@@
IV-B Implementation Details
In this paper, we implement the proposed methods using pytorch and conduct experiments using two NVIDIA RTX 3090 GPUs. For all the experiments, we used the AdamW optimizer and one-cycle learning rate schedule, as well as the same data augmentation strategies. Specifically, we preprocess the training pairs by applying the saturation transform and randomly cropping them to ensure a consistent image size ($320\times 512$ for ETH3D, $384\times 1024$ for Middbury 2014, and $320\times 672$ for other datasets). Below, we provide a detailed description of the training settings for SR-Stereo and DAPE, respectively.
IV-B1 SR-Stereo
IV-B2 DAPE
For the experiments related to the edge estimator, we jointly train the stereo model and the proposed edge estimator on SceneFlow with a batch size of 4 for 50k steps, using a one-cycle learning rate schedule with a learning rate of 0.0001. We use the pre-trained stereo model and edge estimator to generate edge pseudo-labels for target domains. For different datasets, we adopted different settings for the fine-tuning process. For the KITTI, we adopt a batch size of 4 and fine-tune the model for 50k steps with an initial learning rate of 0.0001. As for the ETH3D, we use a batch size of 2 and fine-tune the model for 2,000 steps, also with an initial learning rate of 0.0001. In the case of the Middlebury 2014, we utilize a batch size of 2 and fine-tune the model for 4,000 steps, starting with an initial learning rate of 0.00002.
IV-C Ablation Study
@@ IV-C1 Disparity Stepwise Regression
We explore the optimal settings for DSR as well as its effectiveness. Table I shows the results of DSR in different configurations. In the majority of configurations, the incorporation of DSR (Disparity Stepwise Regression) significantly enhances the performance of the baseline model across different datasets.
@@
Furthermore, we investigate the impact of the number of stepwise regression units employed in the architecture. Experimental results reveal that increasing the number of stepwise regression units leads to better generalization performance of the model.
@@ IV-C3 Number of Stepwise Regression Units
TABLE IV: Extension results for the stepwise regression architecture. For inference, IGEV-Stereo runs 15 disparity updates, while RAFT-Stereo runs 32 disparity updates. RTS:Regression Target Segmentation. Gray: performance is improved after using the proposed method. DCB: Disparity Clip-Balanced Weight ($h$=0.5). Bold: Best.
Methods $m$ SceneFlow Middlebury-H ETH3D
$>3px(\%)$ $>2px(\%)$ $>1px(\%)$
IGEV-Stereo - 3.65 8.44 4.49
IGEV.+RTS 2 3.50 8.09 4.40
3 3.48 8.06 4.71
4 3.48 8.10 4.37
IGEV.+RTS+DCB 2 3.32 7.68 3.87
3 3.26 7.61 4.22
4 3.24 7.78 3.47
RAFT-Stereo - 3.96 14.71 4.51
RAFT.+RTS 3 4.25 15.78 3.86
4 4.01 14.42 4.02
5 3.76 12.26 4.13
RAFT.+RTS+DCB 3 3.70 13.97 3.60
4 3.56 12.64 3.63
5 3.42 11.88 3.59
@@ IV-D Extension of the stepwise regression architecture
To further demonstrate the superiority of the stepwise architecture for disparity estimation and cross-domain generalization, we retain the existing update unit ConvCRU and apply the proposed Regression Target Segmentation and Disparity Clip-Balanced Weight to the iteration-based methods RAFT-Stereo and IGEV-Stereo.
TABLE V: Quantitative evaluation on SceneFlow and KITTI. Our SR-Stereo run 32 disparity updates during inference. Bold: Best.
@@
As shown in Table IV, the proposed method can improve the performance of existing iteration-based methods without changing the network structure. In addition, we can adjust the disparity clip range $m$ to achieve better results according to the characteristics of the existing methods. For instance, in the case of RAFT-Stereo, where the initial disparity is set to 0, a large clip range is required to accelerate the disparity convergence. Therefore, the best results are obtained with a clip range of $m$=5. Conversely, IGEV-Stereo obtains a coarse initial disparity through 3D-CNN, making a smaller clip range more suitable to achieve stable performance improvement.
IV-E Comparison with SOTA Methods
IV-E1 Benchmark Results
In this section, we compare SR-Stereo with the state-of-the-art methods published on SceneFlow and KITTI. As described in section IV-B1, we train the final version of SR-Stereo on SceneFlow. For the KITTI model, we fine-tune the final SceneFlow version of SR-Stereo for 50k steps. Tables V shows the quantitative results. With similar training strategies, SR-Stereo achieves a new SOTA EPE on the SceneFlow test set. Evaluation results on the KITTI benchmark show that SR-Stereo achieves the best performance on the vast majority of metrics. Fig. 7 shows a comparison of the qualitative results of SR-Stereo and IGEV-Stereo on KITTI 2015. Our method is more accurate for edge regions and backgrounds.
IV-E2 Zero-shot Generalization
We test the final SceneFlow version of SR-Stereo directly on Middlebury 2014 and ETH3D. As shown in Table VI, Our SR-Stereo achieves very competitive generalisation performance. Compared with the SOTA method IGEV-Stereo, our method achieves an overall improvement. The qualitative results are shown in Fig 8. As can be seen from the overall range of the disparity maps, our results are closer to the ground truth and show stronger generalization performance.
@@ IV-E2 Zero-shot Generalization
IV-F Domain Adaptation Based on Pre-trained Edge
As mentioned in Section III-A, in addition to proposing a robust network SR-Stereo, we also propose the Domain Adaptation Based on Pre-trained Edge (DAPE) to enhance the performance of existing models after being fine-tuned with sparse ground truth. In this section, we comprehensively evaluate the individual steps of DAPE and demonstrate its effectiveness through experimental results on KITTI, Middbury and ETH3D.
@@ IV-F2 Threshold for Edge Pseudo-label Generation
Results on KITTI: Considering the limitation of the number of KITTI online leaderboard submissions, we have divided the KITTI training set into two parts, with a ratio of 4:1 for fine-tuning the model and evaluation, respectively. The dataset provides sparse ground truth disparities obtained from lidar measurements. Notably, he upper regions of KITTI images primarily consist of sky and distant objects, where ground truth disparities are not available. Moreover, lidar performs poorly in the edge regions of objects, resulting in a lack of ground truth disparities for pixels in these areas. These factors lead to a mismatch between the accuracy ranking and the visualization on the KITTI online leaderboard. To ensure a more objective evaluation of the effectiveness of the proposed DAPE, we present the experimental results of DAPE on different models from both quantitative performance and visualization.
@@
Results on Middbury: We used the additional 13 image pairs provided by Middlebury 2014 to fine-tune the models, and evaluated the domain adaptation performance using the original 15 training image pairs. Similar to the experiments conducted on ETH3D, we employed sparsified ground truth during the fine-tuning process, while relying on the original ground truth for evaluation, as shown in Figure 11.
To verify the performance of DAPE at different resolutions, we evaluated it on Middlebury 2014 with three different resolutions. The qualitative results of DAPE on Middlebury 2014 are presented in Table IX. Consistent with the conclusions from the previous two datasets, the implementation of DAPE with most threshold values leads to performance improvements of the three models. This outcome further confirms the generalizability and robustness of DAPE across different scenarios.
@@ V Conclusion
In this paper, we propose SR-Stereo, a novel stereo matching architecture. Our approach addresses the issue of distribution discrepancies across different datasets by predicting disparity clips. To enhance the accuracy of these disparity clips, we introduce a loss weight that is determined by the regression target scale. Through extensive evaluations on datasets including SceneFlow, KITTI, Middlebury 2014, and ETH3D, we demonstrate that SR-Stereo achieves competitive performance in disparity estimation and state-of-the-art cross-domain generalisation performance. Moreover, we demonstrate that this stepwise regression architecture can be generalised to existing iteration-based methods to improve performance without changing the network structure.
Additionally, we propose a method called Domain Adaptation Based on Pre-trained Edges (DAPE) to address the issue of edge blurring in fine-tuned models with sparse ground truth. Specifically, we use the RGB image and the disparity predicted by the pre-trained stereo matching model to estimate the edge map of the target domain image. This eliminates the need for a complex edge estimation network. The generated edge map is then filtered to generate edge map background pseudo-labels. These pseudo-labels, along with the sparse ground truth disparity on the target domain, are used as supervision to jointly fine-tune the pre-trained stereo matching model. Experimental results at KITTI, Middbury 2014, and ETH3D show that DAPE significantly improves the disparity estimation performance of the fine-tuned model, especially in untextured and detailed regions.
@@
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.

### v3

@@ III-B5 Disparity Clip-Balanced Weight
In SR-Stereo, the regression disparity is divided into multiple clips with the same range to mitigate the distributional discrepancies among different domains. Ideally, a regression disparity of size $L$ is split into a number of disparity clips of size $1.5m$ and a small disparity clip of size $n$ as follows:
@@ III-B6 Loss Function
We use Smooth $L1$ loss to supervise the initial disparity and the disparity clips $d_{k}$ because it is more sensitive to large errors:
@@ III-B7 Generalization of the stepwise regression architecture
We observe that although the core idea is different, the stepwise regression architecture is similar to existing iteration-based methods in terms of processing flow, and thus can be easily generalized based on existing methods. In Section IV-D, we also apply the proposed stepwise regression architecture to the classic iteration-based method RAFT-Stereo, and the experimental results also show that the stepwise regression architecture performs better.
@@ III-C1 Edge Estimator
For the proposed DAPE, the reliability of edge pseudo-label is vital to the performance of the fine-tuned model. Existing edge estimation methods usually take only an RGB image as input and use a series of complex 2D CNNs to estimate the contour of object by extracting high-level features. However, this approach significantly increases the computational cost and limits the generalization performance due to the large discrepancies in object classes among different domains.
Figure 5: Architecture of the edge estimator. The entire architecture consists of only one residual block and one convolution block.
In this paper, we claim that using disparity for edge estimation is a better and more robust approach because disparity estimation focuses on local geometric information, whose representations are similar across domains. Moreover, compared to RGB image, disparity contains almost no object texture information, which greatly reduces the difficulty of edge estimation. Thus, a lightweight edge estimator is proposed to achieve accurate and robust performance by introducing disparity. The architecture of the edge estimator is shown in Figure 5. First, a residual block is used to extract the edge feature in the disparity:
@@ III-C2 Edge Map Background Pseudo-label Generation for Target Domain
The proposed edge estimator achieves edge estimation by introducing disparity, and therefore is affected by the accuracy of the disparity. As shown in Figure 6, existing disparity estimation methods perform badly in reflection regions, which leads to erroneous edges. In this paper, an edge map background pseudo-label is proposed to generate reliable supervision for fine-tuning. Intuitively, in an ideal edge map, both the edge region and non-edge region can be leveraged as supervision to improve the performance of the fine-tuned disparity estimation model. For instance, the disparity map typically exhibits high gradients in edge region, while it remains relatively flat in non-edge region. Erroneous edges in reflection regions do not affect the reliability of the pixels in the non-edge region. Therefore, we can simply utilize pixels in non-edge region as the pseudo-label (i.e., the edge map background pseudo-label):
@@ IV-A Datasets and Evaluation Metrics
Middlebury 2014 [30] consists of two batches of indoor image pairs. The first batch provides 15 training pairs and 15 test pairs, with each scene provided in three different resolutions. The second batch provides 13 additional training pairs, but contains only one resolution. All training pairs are provided with dense hand-labeled ground truth disparity, which ranges from 0 to 300. We directly use the training pairs from the first batch to evaluate the generalization performance of SR-Stereo. For DAPE, we utilize the additional 13 image pairs to fine-tune the model and evaluate its performance using the 15 training pairs from the first batch. The evaluation metric used is the 2-pixel error.
ETH3D [50] is a grayscale image dataset that includes a variety of indoor and outdoor scenes. It consists of 27 training pairs and 20 test pairs. The dataset provides sparsely labeled ground truth disparities for the training pairs, ranging from 0 to 60 (the smallest among several datasets). Similar to the usage in Middbury 2014, we employ the training pairs to directly assess the generalization performance of the proposed method, using the 1-pixel error as the evaluation metric. Additionally, we use 14 training pairs as fine-tuning samples and 13 training pairs as test samples to evaluate the effectiveness of the proposed Domain Adaptation Based on Pre-trained Edge.
@@
IV-B Implementation Details
In this paper, we implement the proposed methods using pytorch and conduct experiments using two NVIDIA RTX 3090 GPUs. For all the experiments, we use the AdamW optimizer and one-cycle learning rate schedule, as well as the same data augmentation strategies. Specifically, we preprocess the training pairs by applying the saturation transform and randomly cropping them to ensure a consistent image size ($320\times 512$ for ETH3D, $384\times 1024$ for Middbury 2014, and $320\times 672$ for other datasets). Below, we provide a detailed description of the training settings for SR-Stereo and DAPE, respectively.
IV-B1 SR-Stereo
IV-B2 DAPE
For the experiments related to the edge estimator, we jointly train the stereo model and the proposed edge estimator on SceneFlow with a batch size of 4 for 50k steps, using a one-cycle learning rate schedule with a learning rate of 0.0001. We use the pre-trained stereo model and edge estimator to generate edge pseudo-labels for target domains. Following existing methods [16, 22, 15], we adopte different settings of fine-tuning process for different datasets. For the KITTI, we adopt a batch size of 4 and fine-tune the model for 50k steps with an initial learning rate of 0.0001. As for the ETH3D, we use a batch size of 2 and fine-tune the model for 2,000 steps, also with an initial learning rate of 0.0001. In the case of the Middlebury 2014, we utilize a batch size of 2 and fine-tune the model for 4,000 steps, starting with an initial learning rate of 0.00002.
IV-C Ablation Study
@@ IV-C1 Disparity Stepwise Regression
We explore the optimal settings for DSR as well as its effectiveness. Table I shows the results of DSR in different configurations. In the majority of configurations, the incorporation of DSR significantly enhances the performance of the baseline model across different datasets.
@@
Furthermore, we investigate the impact of the number of stepwise regression units employed in the architecture. Experimental results reveal that increasing the usage times of stepwise regression units leads to better generalization performance of the model.
@@ IV-D Extension of the stepwise regression architecture
To further demonstrate the superiority of the stepwise architecture for disparity estimation and cross-domain generalization, we retain the existing update unit ConvCRU and apply the proposed Regression Objective Segmentation and Disparity Clip-Balanced Weight to the iteration-based methods RAFT-Stereo and IGEV-Stereo.
TABLE IV: Quantitative evaluation on SceneFlow and KITTI. Our SR-Stereo run 32 disparity updates during inference. Bold: Best.
@@
TABLE V: Extension results for the stepwise regression architecture. For inference, IGEV-Stereo runs 15 disparity updates, while RAFT-Stereo runs 32 disparity updates. ROS:Regression Objective Segmentation. DCB: Disparity Clip-Balanced Weight ($h$=0.5). Gray: performance is improved after using the proposed method. Bold: Best.
Methods $m$ SceneFlow Middlebury-H ETH3D
$>3px(\%)$ $>2px(\%)$ $>1px(\%)$
IGEV-Stereo - 3.65 8.44 4.49
IGEV.+ROS 2 3.50 8.09 4.40
3 3.48 8.06 4.71
4 3.48 8.10 4.37
IGEV.+ROS+DCB 2 3.32 7.68 3.87
3 3.26 7.61 4.22
4 3.24 7.78 3.47
RAFT-Stereo - 3.96 14.71 4.51
RAFT.+ROS 3 4.25 15.78 3.86
4 4.01 14.42 4.02
5 3.76 12.26 4.13
RAFT.+ROS+DCB 3 3.70 13.97 3.60
4 3.56 12.64 3.63
5 3.42 11.88 3.59
As shown in Table V, the proposed method can improve the performance of existing iteration-based methods without changing the network structure. In addition, we can adjust the disparity clip range $m$ to achieve better results according to the characteristics of the existing methods. For instance, in the case of RAFT-Stereo, where the initial disparity is set to 0, a large clip range is required to accelerate the disparity convergence. Therefore, the best results are obtained with a clip range of $m$=5. Conversely, IGEV-Stereo obtains a coarse initial disparity through 3D-CNN, making a smaller clip range more suitable to achieve stable performance improvement.
IV-E Comparison with SOTA Methods
IV-E1 Benchmark Results
In this section, we compare SR-Stereo with the state-of-the-art methods published on SceneFlow and KITTI. As described in Section IV-B1, we train the final version of SR-Stereo on SceneFlow. For the KITTI model, we fine-tune the final SceneFlow version of SR-Stereo for 50k steps. Tables IV shows the quantitative results. With similar training strategies, SR-Stereo achieves a new SOTA EPE on the SceneFlow test set. Evaluation results on the KITTI benchmark show that SR-Stereo achieves the best performance on the vast majority of metrics. Figure 7 shows a comparison of the qualitative results of SR-Stereo and IGEV-Stereo on KITTI 2015. Our method is more accurate for edge regions and backgrounds.
IV-E2 Zero-shot Generalization
We test the final SceneFlow version of SR-Stereo directly on Middlebury 2014 and ETH3D. As shown in Table VI, Our SR-Stereo achieves very competitive generalisation performance. Compared with the SOTA method IGEV-Stereo, our method achieves an overall improvement. The qualitative results are shown in Figure 8. As can be seen from the overall range of the disparity maps, our results are closer to the ground truth and show stronger generalization performance.
@@ IV-E2 Zero-shot Generalization
IV-F Domain Adaptation Based on Pre-trained Edge
As mentioned in Section III-A, in addition to proposing a robust network SR-Stereo, we also propose the DAPE to enhance the performance of existing models after being fine-tuned with sparse ground truth. In this section, we comprehensively evaluate the individual steps of DAPE and demonstrate its effectiveness through experimental results on KITTI, Middbury and ETH3D.
@@ IV-F2 Threshold for Edge Pseudo-label Generation
Results on KITTI: Considering the limitation of the number of KITTI online leaderboard submissions, we have divided the KITTI training set into two parts, with a ratio of 4:1 for fine-tuning the model and evaluation, respectively. The dataset provides sparse ground truth disparities obtained from lidar measurements. Notably, the upper regions of KITTI images primarily consist of sky and distant objects, where ground truth disparities are not available. Moreover, lidar performs poorly in the edge regions of objects, resulting in a lack of ground truth disparities for pixels in these areas. These factors lead to a mismatch between the accuracy ranking and the visualization on the KITTI online leaderboard. To ensure a more objective evaluation of the effectiveness of the proposed DAPE, we present the experimental results of DAPE on different models from both quantitative performance and visualization.
@@
Results on Middbury: We use the additional 13 image pairs provided by Middlebury 2014 to fine-tune the models, and evaluate the domain adaptation performance using the original 15 training image pairs. Similar to the experiments conducted on ETH3D, we employ sparsified ground truth during the fine-tuning process, while relying on the original ground truth for evaluation, as shown in Figure 11.
To verify the performance of DAPE at different resolutions, we evaluate it on Middlebury 2014 with three different resolutions. The qualitative results of DAPE on Middlebury 2014 are presented in Table IX. Consistent with the conclusions from the previous two datasets, the implementation of DAPE with most threshold values leads to performance improvements of the three models. This outcome further confirms the generalizability and robustness of DAPE across different scenarios.
@@ V Conclusion
In this paper, we propose SR-Stereo, a novel stereo matching architecture. Our approach addresses the issue of distribution discrepancies across different datasets by predicting disparity clips. To enhance the accuracy of these disparity clips, we introduce a loss weight that is determined by the regression objective scale. Through extensive evaluations on datasets including SceneFlow, KITTI, Middlebury 2014, and ETH3D, we demonstrate that SR-Stereo achieves competitive performance in disparity estimation and state-of-the-art cross-domain generalisation performance. Moreover, we demonstrate that this stepwise regression architecture can be generalised to existing iteration-based methods to improve performance without changing the network structure.
Additionally, we propose a method called Domain Adaptation Based on Pre-trained Edge (DAPE) to address the issue of edge blurring in fine-tuned models with sparse ground truth. Specifically, we use the RGB image and the disparity predicted by the pre-trained stereo matching model to estimate the edge map of the target domain image. This eliminates the need for a complex edge estimation network. The generated edge map is then filtered to generate edge map background pseudo-labels. These pseudo-labels, along with the sparse ground truth disparity on the target domain, are used as supervision to jointly fine-tune the pre-trained stereo matching model. Experimental results at KITTI, Middbury 2014, and ETH3D show that DAPE significantly improves the disparity estimation performance of the fine-tuned model, especially in untextured and detailed regions.